OpenStudy (anonymous):

what is the equation of the line that passes through (-2,3) and is parallel to 2x+3y=6?

OpenStudy (ash2326):

let's find the solution of the given line 2x+3y=6 3y=-2x+6 y=-2/3x+2 slope of this line =-2/3 line parallel to it will also have slope =-2/3 let that line be y=-2/3x+c it passes through -2,3 so this satisfy the equation 3=+2/3*2+c 3-4/3=c 5/3=c so equation is y=-2/3x+5/3 3y+2x=5

OpenStudy (anonymous):

The equation of a line is y=1/4x-2. What is the equation of the line that is parallel to the first line and passes through (4,-2)?

OpenStudy (mertsj):

The slope of y = 1/4x -2 is 1/4 the parallel line has the same slope. So write the equation of a line whose slope is 1/4 and contains (4,-2)